40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13731     Entry time: Thu Apr 5 13:46:42 2018
Author: gautam 
Type: Update 
Category: SUS 
Subject: big earthquake 

Seems like there was a 5.3 magnitude EQ ~10km from us (though I didn't feel it). All watchdogs were tripped so our mirrors definitely felt it. ITMX is stuck (but all the other optics are damping fine). I tried the usual jiggling of DC bias voltage but ITMX still seems stuck. Probably a good sign that the magnet hasn't come off, but not ideal that I can't shake it free..

edit: after a bit more vigorous shaking, ITMX was freed. I had to move the bias slider by +/-10,000 cts, whereas initially I was trying +/-2000 cts. There is a tendency for the optic to get stuck again once it has been freed (while the optic's free swinging motion damps out), so I had to keep an eye out and as soon as the optic was freed, I re-engaged the damping servos to damp out the optic motion quickly.

Attachment 1: EQ_April52018.png  339 kB  Uploaded Thu Apr 5 14:49:05 2018  | Show | Show all
ELOG V3.1.3-